OpenStudy (anonymous):

Factor The Quadratic Equation : 8x^2 -15x-1=-7x .

OpenStudy (anonymous):

(x-1)(8x+1) = 0 x =1 or, x = -1/8

OpenStudy (anonymous):

mimi cross check ur derived one 8x^2 -6x -1 = 0 7 -15 = -8

OpenStudy (anonymous):

I Am Using The Quadratic Formula And I Still Don't Understand . That's Why I'm Asking For Help .

OpenStudy (mimi_x3):

woops, typo sorry 8x^2 -8x -1=0

OpenStudy (mimi_x3):

Don't understand it ? a = 8 , b=-8 , c = -1 Then sub it into the quadratic formula.

OpenStudy (anonymous):

I Did That . I Got x=8+-4SQRT6 Over 16 .